# Towards Understanding the Fragility of Multilingual LLMs against Fine-Tuning Attacks

### v1

@@ 6 Discussion and Future Work
Our work is the first to reveal that fine-tuning attacks can generalize cross-lingually, where models that are aligned for multilingual safety can be jailbroken through fine-tuning attack in one language. We also identify the language-agnostic parameters within multilingual LLMs that is responsible for safety refusal. Future work on defending LLMs against fine-tuning attacks should robustify this parameters to make multilingual LLMs safer—to the best of our knowledge, all existing work has only focused on English (Hsu et al., 2024; Tamirisa et al., 2024; Huang et al., 2024).
@@
Our work only focuses on cross-lingual generalization of one type of jailbreaking method, which is fine-tuning on harmful datasets. The language coverage of our work is also limited by that of our safety evaluation datasets and safety evaluators. Furthermore, our interpretability experiments, which reveal the language-agnostic safety parameters, focuses on understanding why fine-tuning attack can serve as cross-lingual attack vector. We hope future work can extend our findings to design more robust safety guardrails that are resistant to cross-lingual fine-tuning attacks and make multilingual LLMs safer.
@@
Our research contributes to the responsible development of LLMs by revealing their potential vulnerabilities: fine-tuning attacks can generalize cross-lingually. While we acknowledge that malicious actors exploit cross-lingual transfer of supervised fine-tuning with harmful data to undo safety alignment training that has been conducted in many languages, we believe that identifying the issues is the first critical step to address them. Our findings also suggests that harmful data filtering before fine-tuning for all languages is necessary to mitigate fine-tuning attacks. Our proposed safety information localization method and shared information ratio metric can also better quantify the risks of the cross-lingual transfer of fine-tuning attacks.
@@ Appendix A Fine-tuning attacks details
In this section we go through the specific hyper-parameters that we optimized for attacking the safety aligned LLMs while using the least amount of resources. The attacks are performed as SFT procedures, using data that contain harmful content.

### v2

@@ 6 Discussion and Future Work
Our work is the first to reveal that fine-tuning attacks can generalize cross-lingually, where models that are aligned for multilingual safety can be jailbroken through fine-tuning attack in one language. We also identify the language-agnostic parameters within multilingual LLMs that is responsible for safety refusal. Future work on defending LLMs against fine-tuning attacks should robustify this parameters to make multilingual LLMs safer—to the best of our knowledge, all existing work has only focused on English (Hsu et al., 2024; Tamirisa et al., 2024; Huang et al., 2024). It is also worth exploring whether such findings hold for multimodal LLM safety (Chi et al., 2024).
@@
This work only focuses on the cross-lingual generalization of one type of jailbreaking method, namely fine-tuning on harmful datasets. The language coverage of our work is also limited by that of our safety evaluation datasets and safety evaluators. Furthermore, our interpretability experiments, which reveal the language-agnostic safety parameters, focus on understanding why fine-tuning attacks can serve as cross-lingual attack vectors.
While our study provides important insights into the mechanisms underlying these vulnerabilities, it does not account for other possible attack vectors, such as adversarial prompting or reinforcement learning-based jailbreaks, which may also exhibit cross-lingual transferability. Additionally, our proposed safety information localization method and shared information ratio metric, while useful for assessing risks, require further validation across a wider range of model architectures and multilingual settings.
We hope that future work can extend our findings to design more robust safety guardrails that are resistant to cross-lingual fine-tuning attacks and contribute to making multilingual LLMs safer.
@@
Our research contributes to the responsible development of LLMs by revealing their potential vulnerabilities: fine-tuning attacks can generalize cross-lingually. While we acknowledge that malicious actors exploit cross-lingual transfer of supervised fine-tuning with harmful data to undo safety alignment training that has been conducted in many languages, we believe that identifying the issues is the first critical step to address them. Our findings also suggest that harmful data filtering before fine-tuning for all languages is necessary to mitigate fine-tuning attacks. Our proposed safety information localization method and shared information ratio metric can also better quantify the risks of the cross-lingual transfer of fine-tuning attacks.
Acknowledgments
We thank anonymous reviewers for their constructive suggestions and fruitful discussion. We also extend our sincere gratitude to Diego Garcia-Olano for his early review of this work and for the insightful exchanges that contributed to its development. YH and HZ are partially supported by an NSF IIS grant No. 2416897.
@@ Appendix A Fine-tuning attacks details
In this section, we go through the specific hyper-parameters that we optimized for attacking the safety aligned LLMs while using the least amount of resources. The attacks are performed as SFT procedures, using data that contain harmful content.